OpenStudy (anonymous):

of people on a jury of 12 with German ancestry. You may assume that X is binomially distributed with parameters n=12 and p=0.17. (a) Find the probability that all jury members have German ancestry.

OpenStudy (kropot72):

The question is incomplete, however assuming that the probability that a juror has German ancestry is 0.17 the required probability is \[P(all\ jurors\ have\ German\ ancestry)=\left(\begin{matrix}12 \\ 12\end{matrix}\right)\times(0.17)^{12}\ \times(1-0.17)^{0}=0.17^{12}\ \times0.83\]

OpenStudy (kropot72):

The required probability is \[0.17^{12}\ \times0.83=you\ can\ calculate\]